Tohban Report 2017-02-22

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 8 February 2017
End Date: 15 February 2017
Tohban: Nicole Duncan
Tohban email:
Next Tohban: Pascal
List all reports


Solar Activity

This week was quiet. There is a somewhat interesting AR that has rotated on disk recently, one to keep an eye on, its just popped off a C4 this morning and has a 25% forecast for another C-class event. There are currently 3 names ARs on disk.

How many GOES flares occurred?

 Flares above B, C, M, X class were     14     0     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      2     0     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      2     0     0     0

There were RHESSI flares/GOES flares 10 / 14 over the time range 14-Feb-17 21-Feb-17

Memory Management

Decimation was normal/vigorous.

Last week, on Feb 15 the SSR jumped from ~30% to ~60%. On Friday we rolled back the pointer and erased data for a few days over a non-flaring period. The SSR reduced to ~30% where is has remained after passes for the remainder of the week.

BGS station is still in the red for downlink. Still unclear what the issue is.

Spacecraft Status

D3 and D6 are on. Cold tip temperatures are 128.6 and 127.3 K. Cryocooler power is still at 75W.

The background flux is up at 12%, and the out shutter is set to ~15%. We still have margin, but Ops suggest that we adjust this. The idea is that we will raise the thin shutter out threshold to 5% above the current deadtime, so we will raise it to ~17%.

Data Gaps

Due to a rolled back pointer we have a large data gap for the following period: Feb 13th ~11:30am - Feb 15th ~11:30am

7-048-20:54:20 start sc_ssr_play halted
17-048-20:54:28 start sc_ssr_pnt halted 
17-048-20:55:11 start sc_ssr_play halted
17-048-20:56:27 start tcreset_ssr halted
17-048-20:56:55 start sc_ssr_poff halted
17-048-20:57:04 start sc_ssr_pnt halted
17-048-20:57:20 start sc_ssr_play halted
17-048-20:57:26 start write_ssr_info
17-048-20:57:40 start show_temps_avg
17-048-20:57:57 CP1: 151.4 K  
17-048-20:57:57 CP2: 150.0 K  
17-048-20:57:57 CT1: 126.7 K  
17-048-20:57:57 CT2: 125.0 K  
17-048-20:57:57 TST: 187.8 K  
17-048-20:57:57 cp1 - ct1 = 24.65 K
17-048-20:57:57 Cryo Power: 75.4 W
17-048-20:57:57 Cryo Accel: 12.0 mG

Detector issues

D3 rear continues to have issues with Livetime dropping to 0% at times. Since we are maxed out on the threshold the only knob we have left to turn is the HV, and for the sake of keeping the front segments operating well we decided to keep the HV at the present level.

D6 front fast threshold was increased from 0x80 to 0x90 on Tuesday Feb 21

17-052-17:21:46 start idib_chg_thrshld(6,FRONT,FAST,0x90) ; was 0x80
17-052-17:22:22 start show_temps_avg
17-052-17:22:40 CP1: 152.6 K
17-052-17:22:40 CP2: 151.2 K
17-052-17:22:40 CT1: 127.9 K
17-052-17:22:40 CT2: 126.1 K
17-052-17:22:40 TST: 189.0 K
17-052-17:22:40 cp1 - ct1 = 24.61 K
17-052-17:22:40 Cryo Power: 75.4 W
17-052-17:22:40 Cryo Accel: 12.1 mG


Spacecraft Management

Decimation Normal/Vigorous
HLAT Decimation Rear decimation
Night time data (fronts) plus/minus 4 minutes from daylight
Night time data (rears) plus/minus 4 minutes from daylight
Require extra passes? No
Requirement for moving pointer? Yes
Attenuator operation Commanded out
Detector problems? See notes above.
Personal tools